# multivariateGPT: a decoder-only transformer for multivariate categorical and numeric data

### v1

@@ 1 Introduction
Data collected in complex systems are often a mixture of categorical and numeric values that are informatively and sparsely sampled across a range of classes. Each of these features presents technical challenges, and current approaches are often limited in modeling one or more of these aspects Shukla and Marlin (2020).
@@
One class of models are autoregressive neural network-based sequence models, such as recurrent neural networks, long-short term memory networks Hochreiter and Schmidhuber (1997), and transformers Vaswani et al. (2017). However, these models do not natively handle categorical data (including as text) and numeric data within a single model.
@@
Using language models as a starting point, four common approaches to create a multimodal model that processes numeric and categorical data simultaneously are (a) tokenization of strings representing numeric quantities Singh and Strouse (2024); Gruver et al. (2023) (b) alternate embedding strategy for numeric tokens such as XVal or MMD Zausinger et al. (2024); Alberts et al. (2024)(c) discretization of numeric values Zhu et al. (2024) and (d) shared embedding space with a model designed for the specific type of numeric data used, such as with image data Sun et al. (2023). The first three approaches operate within the existing language modeling framework. However, numeric data from different classes shares a common representation and rely on the context of the preceding token to differentiate meaning. Approaches (a) and (c) both rely on categorical representation of numeric quantities preventing shared information across related tokens: if tokens representing values of $21$ and $23$ are seen in training, there is no information shared with the token representing $22$. Approach (d) introduces additional complexity of training a separate model for the numeric domain or having domain specific encoding and decoding methods. Additional challenges occur when numeric data modalities can have variable size, often requiring resizing or alternative methods before encoding Han et al. (2022b); Tang et al. (2025).
@@
A second class of models are based on neural ordinary differential equations which use a continuous representation of the timeseries Chen et al. (2018). These models handle irregular sampling but cannot represent crossing trajectories. Augmentation addresses this issue Rubanova et al. (2019), but cannot handle stochasticity without conversion to a stochastic differential equation approach which increases complexityKidger et al. (2021). Trajectory flow matching has improved this through simulation-free trainingZhang et al. (2024). However, these models are limited in modeling categorical data and extracting information from sampling measurement timing.
@@
A concrete example of the critical importance of models with these capabilities is in healthcare. Data includes categorical values such as diagnostic codes or text as well as numeric data such as laboratory values, vital signs, and medication dosages. There are thousands of potential classes of data, but only a few are measured at any given time. Additionally, the timing of observations reflect a decision to record an observation, which itself contains information Getzen et al. (2023).
@@ 2.5 Model Architecture
We use a decoder-only transformer to approximate Eq. 1 by minimizing Eq. 8. The model architecture follows the nanoGPT implementation of a decoder-only transformer Karpathy (2022) with modifications to the embedding scheme and loss function to implement methods described in Section 2. A schematic is shown in Figure 1. In this implementation, a single feed-forward layer is used for value mapping. The final linear layer consists of a Class Head and Value Head which are used to estimate the loss based on the joint probability of the next token’s class and value.
@@ 3.1 Fitting and Generalization in Simple Harmonic Oscillators
Trajectories of damped simple harmonic oscillators provide a toy system that requires the ability to model second order dynamics and crossing trajectories. Neural ordinary differential equation models, canonical conditional flow matching, and bridge matching cannot fit these trajectories due to the presence of crossing trajectories Zhang et al. (2024). We compared the multivariateGPT model to discrete token transformer baselines on two tasks: trajectory reconstruction and generalization.
@@ 3.2 Clinical Data sets
    eICU Sepsis Data: subset of patients in the eICU Collaborative Research Database v2.0 with a diagnosis of sepsis, chosen to match Zhang et al. (2024). Includes 3,362 patients, with data on age, sex, heart rate, mean arterial pressure (MAP), and norepinephrine infusion rate during the first 24 hours of ICU admission. This was derived from the eICU collaborative database Pollard et al. (2018).
@@
    MIMIC-IV Electrocardiogram Data: subset of 89,339 electrocardiograms from the MIMIC-IV electrocardiogram database. Each record includes patient age, sex, and voltages for the 12 standard leads of an electrocardiogram. This was taken from MIMIC-IV-ECG dataset Gow et al. (2023) derived from the MIMIC-IV database Johnson et al. (2023).
@@
    Physionet ICU Data: data on 65,155 patients admitted to the the ICU from 3 hospital systems with demographics and hourly records of vital signs and lab values representing 36 unique classes Moor et al. (2021).
@@
We evaluated the multivariateGPT model compared to discrete-token transformer baselines, an open-source LLM, and TFM-ODE, which recently outperformed a range of neural ODE and SDE methods in this data set. This data set has irregularly and potentially informatively sampled heart rate and MAP measurements. As in Zhang et al. (2024), we focused on a subset of patients admitted with sepsis as their primary diagnosis. Two tasks were tested due to differences in what each model class was able to predict. While the multivariateGPT and discrete-token transformer models predict both measurements and the time between measurements, TFM-ODE requires specifying the time at which measurements occur.
@@ 4 Related Work
Token-Based methods: A naive discrete token-based approach is not optimal for numeric values Spathis and Kawsar (2024). Careful tokenization of strings containing digits can improve performance on tasks Born and Manica (2023), and single digit tokenization has been found to be more efficient than tokenization into longer groups Zhou et al. (2024); Schmidt et al. (2024). The LLAMA series of models adopted single-digit-based tokenization Touvron et al. (2023) and an adaptation of these models for time series modeling enforced digit-based tokenization through addition of delimiting characters Gruver et al. (2023). Right-to-left digit tokenization was shown to improve arithmetic performance over naive byte pair encoding strategies across a range of arithmetic tasks Singh and Strouse (2024). Additional efforts to transform numeric quantities into discrete tokens include discretization into bins Gorishniy et al. (2022); Stein et al. (2024); Rajkomar et al. (2018); Renc et al. (2024); Zhu et al. (2024); Huang et al. and a similarity-driven quantization approach for time series data Zhicheng et al. (2024). This is the main method employed in current models of medical data Rajkomar et al. (2018); Theodorou et al. (2023); Pang et al. (2024); Renc et al. (2024). Within image and audio processing, multiple methods have been developed for transforming signals into discrete tokens including LFQ, FSQ, and BSQ which utilize vector quantization via discrete codebooks Yu et al. (2023); Jia et al. (2025); Mentzer et al. (2023); Zhao et al. (2024). Embedding approaches aligned with the continuous nature of numeric values have also been developed Wang et al. (2025); Golkar et al. (2024). Once tokens are created, standard cross entropy loss is not well suited for numeric decoding and alternate loss functions including regression-like loss Zausinger et al. (2024); Han et al. (2022a) have been developed. A remaining limitation from this existing work is a numeric representation for diverse classes of data with a single likelihood-based loss which we provide here.
Continuous Models Continuous-time trajectory models are an alternative approach to model irregularly sampled data. Models such as the latent trajectory-based Neural-ODEs Rubanova et al. (2019); Chen et al. (2018) and Neural-SDEs Kidger et al. (2021); Oh et al. (2024) have outperformed LSTM- and RNN-based methods Che et al. (2018); Cao et al. (2018). In turn, trajectory flow matching (TFM) outperformed these models in clinical timeseries modeling with improved computational efficiency through its simulation-free training and the additional benefit of estimating uncertainty Zhang et al. (2024). Additional methods to capture uncertainty include Brouwer et al. (2022). These methods handle irregularly spaced inputs, but they do not currently derive information from the irregular spacing and do not handle categorical data well.
@@ A.2 MIMIC-IV Electrocardiogram
For TFM-ODE we used the implementation that was published for this data set in Zhang et al. [2024]. For large-language model baseline, we used gemma 1.1-7b Team et al. [2024]. The following prompt was used: "You are tasked with predicting future values of heart rate and mean arterial pressure. the input data is: <data in csv format> give your best estimate to what hr and map will be at the following time: <times for estimation> and The norepi_inf at these times is: <norepi infusion rates> use csv output with the following columns: time, hr, map."
@@ A.4 Physionet ICU Data
The Physionet ICU data consisted of 65,155 unique records for patients admitted to the ICU. Data include demographics, vital signs, and lab values totaling 36 unique categorical and numeric measurement classes Moor et al. [2021]. Numeric classes were z-score normalized except for SpO2 which used a logistic scaling due to a ceiling effect where values cannot be greater than 100% and are commonly at or near this value.

### v2

@@ 1 Introduction
Data collected in complex systems are often a mixture of categorical and numeric values that are informatively and sparsely sampled across a range of classes. Each of these features presents technical challenges, and current approaches are often limited in modeling one or more of these aspects (Shukla and Marlin, 2020).
@@
One class of models are autoregressive neural network-based sequence models, such as recurrent neural networks, long-short term memory networks (Hochreiter and Schmidhuber, 1997), and transformers (Vaswani et al., 2017). However, these models do not natively handle categorical data (including as text) and numeric data within a single model.
@@
Using language models as a starting point, four common approaches to create a multimodal model that processes numeric and categorical data simultaneously are (a) tokenization of strings representing numeric quantities (Singh and Strouse, 2024; Gruver et al., 2023) (b) alternate embedding strategy for numeric tokens such as XVal or MMD (Zausinger et al., 2024; Alberts et al., 2024)(c) discretization of numeric values (Zhu et al., 2024) and (d) shared embedding space with a model designed for the specific type of numeric data used, such as with image data (Sun et al., 2023). The first three approaches operate within the existing language modeling framework. However, numeric data from different classes shares a common representation and rely on the context of the preceding token to differentiate meaning. Approaches (a) and (c) both rely on categorical representation of numeric quantities preventing shared information across related tokens: if tokens representing values of $21$ and $23$ are seen in training, there is no information shared with the token representing $22$. Approach (d) introduces additional complexity of training a separate model for the numeric domain or having domain specific encoding and decoding methods. Additional challenges occur when numeric data modalities can have variable size, often requiring resizing or alternative methods before encoding (Han et al., 2022b; Tang et al., 2025).
@@
A second class of models are based on neural ordinary differential equations which use a continuous representation of the timeseries (Chen et al., 2018). These models handle irregular sampling but cannot represent crossing trajectories. Augmentation addresses this issue (Rubanova et al., 2019), but cannot handle stochasticity without conversion to a stochastic differential equation approach which increases complexity(Kidger et al., 2021). Trajectory flow matching has improved this through simulation-free training(Zhang et al., 2024). However, these models are limited in modeling categorical data and extracting information from sampling measurement timing.
@@
A concrete example of the critical importance of models with these capabilities is in healthcare. Data includes categorical values such as diagnostic codes or text as well as numeric data such as laboratory values, vital signs, and medication dosages. There are thousands of potential classes of data, but only a few are measured at any given time. Additionally, the timing of observations reflect a decision to record an observation, which itself contains information (Getzen et al., 2023).
@@ 2.5 Model Architecture
We use a decoder-only transformer to approximate Eq. 1 by minimizing Eq. 8. The model architecture follows the nanoGPT implementation of a decoder-only transformer (Karpathy, 2022) with modifications to the embedding scheme and loss function to implement methods described in Section 2. A schematic is shown in Figure 1. In this implementation, a single feed-forward layer is used for value mapping. The final linear layer consists of a Class Head and Value Head which are used to estimate the loss based on the joint probability of the next token’s class and value.
@@ 3.1 Fitting and Generalization in Simple Harmonic Oscillators
Trajectories of damped simple harmonic oscillators provide a toy system that requires the ability to model second order dynamics and crossing trajectories. Neural ordinary differential equation models, canonical conditional flow matching, and bridge matching cannot fit these trajectories due to the presence of crossing trajectories (Zhang et al., 2024). We compared the multivariateGPT model to discrete token transformer baselines on two tasks: trajectory reconstruction and generalization.
@@ 3.2 Clinical Data sets
    eICU Sepsis Data: subset of patients in the eICU Collaborative Research Database v2.0 with a diagnosis of sepsis, chosen to match (Zhang et al., 2024). Includes 3,362 patients, with data on age, sex, heart rate, mean arterial pressure (MAP), and norepinephrine infusion rate during the first 24 hours of ICU admission. This was derived from the eICU collaborative database (Pollard et al., 2018).
@@
    MIMIC-IV Electrocardiogram Data: subset of 89,339 electrocardiograms from the MIMIC-IV electrocardiogram database. Each record includes patient age, sex, and voltages for the 12 standard leads of an electrocardiogram. This was taken from MIMIC-IV-ECG dataset (Gow et al., 2023) derived from the MIMIC-IV database (Johnson et al., 2023).
@@
    Physionet ICU Data: data on 65,155 patients admitted to the the ICU from 3 hospital systems with demographics and hourly records of vital signs and lab values representing 36 unique classes (Moor et al., 2021).
@@
We evaluated the multivariateGPT model compared to discrete-token transformer baselines, an open-source LLM, and TFM-ODE, which recently outperformed a range of neural ODE and SDE methods in this data set. This data set has irregularly and potentially informatively sampled heart rate and MAP measurements. As in (Zhang et al., 2024), we focused on a subset of patients admitted with sepsis as their primary diagnosis. Two tasks were tested due to differences in what each model class was able to predict. While the multivariateGPT and discrete-token transformer models predict both measurements and the time between measurements, TFM-ODE requires specifying the time at which measurements occur.
@@ 4 Related Work
Token-Based methods: A naive discrete token-based approach is not optimal for numeric values (Spathis and Kawsar, 2024). Careful tokenization of strings containing digits can improve performance on tasks (Born and Manica, 2023), and single digit tokenization has been found to be more efficient than tokenization into longer groups (Zhou et al., 2024; Schmidt et al., 2024). The LLAMA series of models adopted single-digit-based tokenization (Touvron et al., 2023) and an adaptation of these models for time series modeling enforced digit-based tokenization through addition of delimiting characters (Gruver et al., 2023). Right-to-left digit tokenization was shown to improve arithmetic performance over naive byte pair encoding strategies across a range of arithmetic tasks (Singh and Strouse, 2024). Additional efforts to transform numeric quantities into discrete tokens include discretization into bins (Gorishniy et al., 2022; Stein et al., 2024; Rajkomar et al., 2018; Renc et al., 2024; Zhu et al., 2024; Huang et al., ) and a similarity-driven quantization approach for time series data (Zhicheng et al., 2024). This is the main method employed in current models of medical data (Rajkomar et al., 2018; Theodorou et al., 2023; Pang et al., 2024; Renc et al., 2024). Within image and audio processing, multiple methods have been developed for transforming signals into discrete tokens including LFQ, FSQ, and BSQ which utilize vector quantization via discrete codebooks (Yu et al., 2023; Jia et al., 2025; Mentzer et al., 2023; Zhao et al., 2024). Embedding approaches aligned with the continuous nature of numeric values have also been developed (Wang et al., 2025; Golkar et al., 2024). Once tokens are created, standard cross entropy loss is not well suited for numeric decoding and alternate loss functions including regression-like loss (Zausinger et al., 2024; Han et al., 2022a) have been developed. A remaining limitation from this existing work is a numeric representation for diverse classes of data with a single likelihood-based loss which we provide here.
Continuous Models Continuous-time trajectory models are an alternative approach to model irregularly sampled data. Models such as the latent trajectory-based Neural-ODEs (Rubanova et al., 2019; Chen et al., 2018) and Neural-SDEs (Kidger et al., 2021; Oh et al., 2024) have outperformed LSTM- and RNN-based methods (Che et al., 2018; Cao et al., 2018). In turn, trajectory flow matching (TFM) outperformed these models in clinical timeseries modeling with improved computational efficiency through its simulation-free training and the additional benefit of estimating uncertainty (Zhang et al., 2024). Additional methods to capture uncertainty include (Brouwer et al., 2022). These methods handle irregularly spaced inputs, but they do not currently derive information from the irregular spacing and do not handle categorical data well.
@@ A.2 MIMIC-IV Electrocardiogram
For TFM-ODE we used the implementation that was published for this data set in [Zhang et al., 2024]. For large-language model baseline, we used gemma 1.1-7b [Team et al., 2024]. The following prompt was used: "You are tasked with predicting future values of heart rate and mean arterial pressure. the input data is: <data in csv format> give your best estimate to what hr and map will be at the following time: <times for estimation> and The norepi_inf at these times is: <norepi infusion rates> use csv output with the following columns: time, hr, map."
@@ A.4 Physionet ICU Data
The Physionet ICU data consisted of 65,155 unique records for patients admitted to the ICU. Data include demographics, vital signs, and lab values totaling 36 unique categorical and numeric measurement classes [Moor et al., 2021]. Numeric classes were z-score normalized except for SpO2 which used a logistic scaling due to a ceiling effect where values cannot be greater than 100% and are commonly at or near this value.